1. Consider the data below:

Date

Nominal GDP (billions)

GDP Deflator

Population Growth Rate

2010

$16,015.40

104.62

2011

$16,670.20

106.12

1.8%

2012

$16,324.50

107.44

1.7%

2013

$16,680.60

108.98

.98%

Compute economic growth rates for the US for 2010-11, 2011-12, and 2012-13.
Answer #1

Answer : Real GDP of 2010 = (Nominal GDP / GDP deflator) *100

=> Real GDP of 2010 = (16,015.40 / 104.62) * 100 = $15,308.16

Real GDP of 2011 = (Nominal GDP / GDP deflator) *100

=> Real GDP of 2011 = (16,670.20 / 106.12) * 100 = $15,708.82

Real GDP of 2012 = (Nominal GDP / GDP deflator) *100

=> Real GDP of 2012 = (16,324.50 / 107.44) * 100 = $15,194.06

Real GDP of 2013 = (Nominal GDP / GDP deflator) *100

=> Real GDP of 2013 = (16,680.60 / 108.98) * 100 = $15,306.11

Economic growth rate 2010-11 = [(Real GDP of 2011 - Real GDP of 2010) / Real GDP of 2010] * 100

=> Economic growth rate 2010-11 = [(15,708.82 - 15,308.16) / 15,308.16] * 100

=> Economic growth rate 2010-11 = 2.62%

Economic growth rate 2011-12 = [(Real GDP of 2012 - Real GDP of 2011) / Real GDP of 2011] * 100

=> Economic growth rate 2011-12 = [(15,194.06 - 15,708.82) / 15,708.82] * 100

=> Economic growth rate 2011-12 = - 3.28%

Economic growth rate 2012-13 = [(Real GDP of 2013 - Real GDP of 2012) / Real GDP of 2012] * 100

=> Economic growth rate 2012-13 = [(15,306.11 - 15,194.06) / 15,194.06] * 100

=>  Economic growth rate 2012-13 = 0.74%